Your younger sibling has saved x dollars. You have saved $50 less than the amount your younger sibling has saved. You saved a total of $150.
A) write an algebraic solution that describes the solution above.

B) describe in words how you would solve this equation

C) who saved more money? Explain
Please Need HELP NOW !!! This problem is worth 12 points on my homework assignment ! Please !

Answer by mananth(16949)

Your younger sibling has saved x dollars. You have saved $50 less than the amount your younger sibling has saved. You saved a total of $150.
A) write an algebraic solution that describes the solution above.

B) describe in words how you would solve this equation

C) who saved more money? Explain
younger sibling -------------$x
you save -------$x-50
Total saved by both of you
x+x-50=150
2x-50=150
2x=150+50
2x=200
/2
x=100 sibling saved 100 dollars
Sibling saved $100 . Total saved is $150
so I saved $50
Sibling saved $50 more than me.